# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ 7 Examples of real-world data sets
The first example data set is the Facebook social network between students in California Institute of Technology (Caltech), collected by Traud et al. (2012). Each node in the network is a student and the edges are the Facebook connections between students. We process the data set following Wang and Rohe (2016) and Li et al. (2022) using the 2-core algorithm. The resulting data set contains 734 students with an average degree of 45.29. Similar to our previous experiments, we randomly select half of the students to hold out. The network of the other half of the students ($n=367$) is then privatized and released. The inner product latent space model is used for model fitting, and the latent dimension $d=6$ is selected by the edge cross-validation method of Li et al. (2020a) on the hold-out network. We enforce a strong privacy requirement at the level of $\varepsilon=1$. The distributions of the five local statistics in the released network are calculated and compared with the distributions in the true network.
Table 4: Node-level statistic distribution preservation results with $n=4000$ under the RDPG model, measured by the average Wasserstein distance over 100 replications. The numbers in the parentheses are the corresponding standard errors.
Metric $d$ $\varepsilon$ $\rho=0.025$ $\rho=0.05$ $\rho=0.1$
Hat GRAND Laplace Hat GRAND Laplace Hat GRAND Laplace
Node Degree 3 1 0.017 (<0.001<0.001) 0.026 (0.001) 2.496 (0.004) 0.009 (<0.001<0.001) 0.019 (0.001) 2.053 (0.002) 0.005 (<0.001<0.001) 0.02 (0.001) 1.516 (0.001)
2 0.017 (<0.001<0.001) 0.025 (0.001) 1.613 (0.006) 0.009 (<0.001<0.001) 0.021 (0.001) 1.43 (0.004) 0.005 (<0.001<0.001) 0.019 (0.001) 1.138 (0.003)
5 0.017 (<0.001<0.001) 0.023 (0.001) 0.544 (0.002) 0.009 (<0.001<0.001) 0.018 (0.001) 0.535 (0.002) 0.005 (<0.001<0.001) 0.016 (0.001) 0.511 (0.002)
10 0.017 (<0.001<0.001) 0.025 (0.001) 0.504 (0.002) 0.009 (<0.001<0.001) 0.016 (0.001) 0.487 (0.002) 0.005 (<0.001<0.001) 0.015 (0.001) 0.473 (0.001)
6 1 0.019 (<0.001<0.001) 0.021 (0.001) 2.908 (0.001) 0.01 (<0.001<0.001) 0.016 (0.001) 2.259 (0.001) 0.005 (<0.001<0.001) 0.013 (0.001) 1.598 (0.001)
2 0.019 (<0.001<0.001) 0.021 (0.001) 2.562 (0.003) 0.01 (<0.001<0.001) 0.015 (0.001) 2.038 (0.002) 0.005 (<0.001<0.001) 0.013 (0.001) 1.471 (0.001)
5 0.019 (<0.001<0.001) 0.022 (0.001) 1.308 (0.005) 0.01 (<0.001<0.001) 0.015 (0.001) 1.074 (0.004) 0.005 (<0.001<0.001) 0.013 (0.001) 0.819 (0.003)
10 0.019 (<0.001<0.001) 0.021 (0.001) 0.523 (0.002) 0.01 (<0.001<0.001) 0.015 (0.001) 0.562 (0.002) 0.005 (<0.001<0.001) 0.011 (0.001) 0.593 (0.002)
V-shape 3 1 0.034 (<0.001<0.001) 0.054 (0.002) 5.025 (0.008) 0.018 (<0.001<0.001) 0.038 (0.002) 4.121 (0.004) 0.01 (<0.001<0.001) 0.04 (0.002) 3.039 (0.003)
2 0.034 (<0.001<0.001) 0.05 (0.002) 3.253 (0.012) 0.018 (<0.001<0.001) 0.042 (0.002) 2.872 (0.007) 0.01 (<0.001<0.001) 0.038 (0.002) 2.283 (0.005)
5 0.034 (<0.001<0.001) 0.048 (0.002) 1.105 (0.005) 0.018 (<0.001<0.001) 0.037 (0.002) 1.08 (0.004) 0.01 (<0.001<0.001) 0.032 (0.001) 1.027 (0.004)
10 0.034 (<0.001<0.001) 0.051 (0.002) 1.06 (0.004) 0.018 (<0.001<0.001) 0.031 (0.001) 1.005 (0.003) 0.01 (<0.001<0.001) 0.031 (0.001) 0.962 (0.003)
6 1 0.039 (<0.001<0.001) 0.043 (0.001) 5.846 (0.002) 0.02 (<0.001<0.001) 0.032 (0.001) 4.532 (0.001) 0.01 (<0.001<0.001) 0.027 (0.001) 3.202 (0.001)
2 0.039 (<0.001<0.001) 0.042 (0.001) 5.153 (0.005) 0.02 (<0.001<0.001) 0.031 (0.001) 4.09 (0.004) 0.01 (<0.001<0.001) 0.026 (0.001) 2.948 (0.002)
5 0.039 (<0.001<0.001) 0.044 (0.001) 2.638 (0.01) 0.02 (<0.001<0.001) 0.031 (0.001) 2.158 (0.008) 0.01 (<0.001<0.001) 0.026 (0.001) 1.643 (0.006)
10 0.039 (<0.001<0.001) 0.043 (0.001) 1.061 (0.004) 0.02 (<0.001<0.001) 0.03 (0.001) 1.134 (0.003) 0.01 (<0.001<0.001) 0.023 (0.001) 1.192 (0.004)
Triangle 3 1 0.037 (<0.001<0.001) 0.063 (0.003) 8.075 (0.008) 0.019 (<0.001<0.001) 0.047 (0.003) 6.532 (0.004) 0.01 (<0.001<0.001) 0.05 (0.003) 4.788 (0.003)
2 0.037 (<0.001<0.001) 0.058 (0.002) 5.977 (0.014) 0.019 (<0.001<0.001) 0.05 (0.003) 5.117 (0.008) 0.01 (<0.001<0.001) 0.048 (0.003) 3.961 (0.006)
5 0.037 (<0.001<0.001) 0.051 (0.002) 2.254 (0.012) 0.019 (<0.001<0.001) 0.042 (0.002) 2.091 (0.009) 0.01 (<0.001<0.001) 0.04 (0.002) 1.845 (0.006)
10 0.037 (<0.001<0.001) 0.051 (0.002) 1.34 (0.007) 0.019 (<0.001<0.001) 0.035 (0.001) 1.385 (0.005) 0.01 (<0.001<0.001) 0.038 (0.002) 1.358 (0.005)
6 1 0.057 (0.001) 0.054 (0.002) 8.954 (0.002) 0.023 (<0.001<0.001) 0.039 (0.002) 6.993 (0.002) 0.011 (<0.001<0.001) 0.034 (0.002) 5.009 (0.002)
2 0.057 (0.001) 0.056 (0.002) 8.162 (0.006) 0.023 (<0.001<0.001) 0.038 (0.002) 6.501 (0.004) 0.011 (<0.001<0.001) 0.034 (0.002) 4.735 (0.003)
5 0.057 (0.001) 0.054 (0.002) 5.015 (0.014) 0.023 (<0.001<0.001) 0.036 (0.001) 4.226 (0.01) 0.011 (<0.001<0.001) 0.033 (0.002) 3.274 (0.007)
10 0.057 (0.001) 0.046 (0.001) 2.152 (0.01) 0.023 (<0.001<0.001) 0.034 (0.001) 2.036 (0.007) 0.011 (<0.001<0.001) 0.028 (0.002) 1.92 (0.006)
Eigen Centrality 3 1 0.035 (0.002) 0.035 (0.002) 0.142 (0.002) 0.02 (0.001) 0.023 (0.002) 0.125 (0.002) 0.013 (0.001) 0.016 (0.001) 0.133 (0.001)
2 0.035 (0.002) 0.034 (0.002) 0.217 (0.001) 0.02 (0.001) 0.023 (0.001) 0.178 (0.001) 0.013 (0.001) 0.015 (0.001) 0.161 (0.001)
5 0.035 (0.002) 0.035 (0.002) 0.329 (0.002) 0.02 (0.001) 0.022 (0.001) 0.299 (0.001) 0.013 (0.001) 0.017 (0.001) 0.214 (0.001)
10 0.035 (0.002) 0.034 (0.002) 0.297 (0.003) 0.02 (0.001) 0.023 (0.002) 0.305 (0.002) 0.013 (0.001) 0.016 (0.001) 0.277 (0.002)
6 1 0.049 (0.002) 0.038 (0.002) 0.247 (0.003) 0.022 (0.001) 0.024 (0.002) 0.234 (0.003) 0.014 (0.001) 0.023 (0.002) 0.213 (0.003)
2 0.049 (0.002) 0.036 (0.002) 0.093 (0.002) 0.022 (0.001) 0.025 (0.002) 0.117 (0.001) 0.014 (0.001) 0.021 (0.002) 0.137 (0.002)
5 0.049 (0.002) 0.04 (0.003) 0.309 (0.002) 0.022 (0.001) 0.028 (0.002) 0.255 (0.002) 0.014 (0.001) 0.022 (0.002) 0.213 (0.001)
10 0.049 (0.002) 0.039 (0.002) 0.379 (0.002) 0.022 (0.001) 0.023 (0.002) 0.355 (0.002) 0.014 (0.001) 0.021 (0.002) 0.287 (0.002)
Harmonic Centrality 3 1 3.557 (0.062) 6.924 (0.415) 638.38 (2.005) 1.162 (0.018) 2.385 (0.11) 659.59 (1.106) 0.844 (0.01) 3.485 (0.202) 666.958 (0.861)
2 3.557 (0.062) 6.257 (0.351) 278.025 (2.062) 1.162 (0.018) 2.558 (0.123) 345.161 (1.525) 0.844 (0.01) 3.31 (0.175) 438.446 (1.367)
5 3.557 (0.062) 5.938 (0.351) 100.041 (1.132) 1.162 (0.018) 2.3 (0.097) 133.62 (0.664) 0.844 (0.01) 2.769 (0.122) 172.796 (0.717)
10 3.557 (0.062) 6.585 (0.336) 138.617 (1.075) 1.162 (0.018) 1.975 (0.076) 124.544 (0.664) 0.844 (0.01) 2.6 (0.116) 114.992 (0.443)
6 1 3.716 (0.052) 4.905 (0.21) 911.974 (0.594) 1.042 (0.011) 1.67 (0.058) 834.276 (0.415) 0.909 (0.012) 2.489 (0.113) 761.823 (0.306)
2 3.716 (0.052) 4.798 (0.252) 671.45 (1.439) 1.042 (0.011) 1.617 (0.057) 666.549 (1.167) 0.909 (0.012) 2.435 (0.114) 659.242 (0.767)
5 3.716 (0.052) 5.245 (0.284) 218.088 (1.249) 1.042 (0.011) 1.596 (0.051) 239.014 (1.211) 0.909 (0.012) 2.404 (0.106) 316.235 (1.292)
10 3.716 (0.052) 5.271 (0.241) 74.745 (0.494) 1.042 (0.011) 1.528 (0.063) 120.293 (0.574) 0.909 (0.012) 2.084 (0.109) 170.016 (0.804)
Figure 2 displays the resulting distributions of the five local statistics in four networks as introduced in Section 6: the original network (True), the network released from our method (GRAND), the network generated using the naive Laplace mechanism (Laplace) and the non-private network from standard model estimation (Hat). It can be seen that the privatized network from our method matches the true network well in all of the five metrics. It also substantially outperforms the naive Laplace method, which completely misses the pattern. Meanwhile, GRAND maintains a slightly deviated but similar performance compared to the non-private Hat network. In particular, the harmonic closeness of the original network exhibits a bi-modal pattern, with one tiny lower mode (on the left end of the figure panel). This subtle pattern is also well captured by our method.
@@
Figure 2: The distributions of five local statistics of the privatized Caltech social network with privacy budget $\varepsilon=1$, compared with those in the true network and the non-private “Hat” network.
Statisticians’ collaboration network
The second example data set is the collaboration network between statistical researchers based on publication data in four statistical journals during the period of 2004–2013, originally collected by Ji and Jin (2016). Here we use the processed version of the data set analyzed in Li et al. (2020b). Each node in this network is a statistician and an edge indicates that the two statisticians coauthored at least one paper during the data collection period. We use the same procedure to process the data set as before, and the resulting network has 509 nodes with an average degree of 4.24. Note that this network is much sparser than the Caltech network, indicating a more difficult model fitting. We still hold out half of the nodes and privatize the network structure of the other half using privacy budget $\varepsilon=1$. The inner product latent space model is used and the latent dimension $d=4$ is selected by cross-validation on the hold-out data.
The distributions of the five local statistics in the resulting networks are shown in Figure 3. The advantage of our method is still evident. The released network matches the true network reasonably well, and maintains a performance that is very close to that of the Hat network, with a small deviation due to the incorporation of privacy guarantees. Similar to the previous Caltech network, the true network also exhibits a bi-modal pattern for the harmonic centrality. However, compared with the previous example, the two modes in this case are much closer to each other, which significantly increases the difficulty in preserving them when perturbation is introduced for privacy. The privatized network from our method, though does not perfectly recover the magnitudes, still captures the bi-modal pattern. We consider this as an impressive success, especially under such a small privacy budget.
@@
Figure 3: The distributions of five local properties of the privatized statisticians’ coauthorship network with privacy budget $\varepsilon=1$, compared with those in the true network and the non-private “Hat” network.
Additional evaluations on social network data
We also evaluate our methods on 107 social networks from the CommunityFitNet data set introduced in Ghasemian et al. (2019). Our method demonstrates competitive performance, and uniformly outperforms the naive Laplace mechanism. For brevity, we include the summary results in Appendix G.
@@ 8 Discussion
This paper defines node-level differential privacy and introduces a novel privatization mechanism, named GRAND, to achieve node-level differential privacy. Our work combines data privacy with a novel node-wise estimation method, yielding a mechanism to release network structures under node-level differential privacy while provably preserving original network properties asymptotically for a general class of network models. Notably, this is the first computationally feasible node DP mechanism for network release that includes guarantees for property preservation. Numerically, we demonstrate its effectiveness in preserving the distributions of multiple local statistics in both synthetic and real-world data sets.
It is important to note that our mechanism offers a flexible framework, where various components can be substituted with alternative options. For instance, the choice of network estimation method for the hold-out network, the criterion for node-wise estimation, and the DIP mechanism for introducing noise can be adapted as appropriate. Thus, while laying a foundational cornerstone for solving the node DP problem, this framework also opens numerous avenues for future research and customization. One promising direction is to generalize the proposed method so that it can encode more systematic privacy protection for the hold-out network. In fact, as previously discussed, using such hold-out data currently appears unavoidable to achieve the desired privacy guarantees within the current scope of generality. One viable solution is to adopt a $(\varepsilon,\delta)$-differentially private holdout set (e.g., through kernel density estimation) with theoretical underpinnings and a transfer learning model as discussed in Section 4.2. Yet another potentially useful future direction is to extend the current framework to accommodate graphon models, in which $Z_{i}$’s are univariate with the known uniform distribution but $W$ is unknown.
@@ Appendix A Proof of Theorem 1
Due the node-wise estimation procedure, note that the change of any node $i\in[n]$ does not change the estimate of $\hat{Z}_{j},j\neq i,j\in[n]$. That means, the estimates $\hat{Z}_{i}$’s are separable. By applying Theorem 2 of Bi and Shen [2023], $\tilde{Z}_{i}$ acquired via Equation (8) is differentially private. Since the generative function $W$ is assumed to be known. We have each node of $\tilde{A}^{11}$ also being differentially private. ∎
@@ Appendix B Proof of Theorem 2 (Marginal Convergence)
The DIP procedure, given $\hat{Z}_{1}$, is to apply univariate DIP transformation by columns. In each step, we use the approximated (conditional) CDF, $\hat{F}_{m}^{j|1:(j-1)}$ of the $j$th variable, conditioning on the previous various j1j-1 variables, for privatization: for each $i$, we generate $e_{i}$ from $\text{Laplace}(0,1/\epsilon)$ independently, and then compute
$\tilde{Z}_{1ij}=(\hat{F}_{m}^{j|1:(j-1)})^{-1}(G(\hat{F}_{m}^{j|1:(j-1)}(\hat{Z}_{1ij})+e_{i})).$
@@ B.1 Crucial tools
Let $Q=\{Q_{i.}\}_{i=1}^{m}$ be i.i.d. random variables from a distribution $F$ on $\bf R$}, and let $F_{m}(x)$ be the empirical CDF computed from $Q$. Let $\hat{Q}=\{\hat{Q}_{i.}\}_{i=1}^{m}$ be a perturbed version of $Q$, which can be potentially dependent. Let $\hat{F}_{m}(x)$ denote the empirical CDF based on the perturbed data $\hat{Q}$.
@@
  • The perturbed data satisfies $\sup_{1\leq i\leq m}|\hat{Q}_{i.}-Q_{i.}|\leq\delta_{m}$ for some $\delta_{m}$ that is associated with $m$.
@@
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\leq O_{P}\left(\delta_{m}\right)+O_{P}\left(\sqrt{\frac{1}{m}}\right).$
@@
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\leq\underbrace{\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F_{m}(x)|}_{\text{Perturbation Error}}+\underbrace{\sup_{x\in{\bf R}}|F_{m}(x)-F(x)|}_{\text{Sampling Error}}.$
From Glivenko-Cantelli, we have:
$\sup_{x\in{\bf R}}|F_{m}(x)-F(x)|=O_{P}\left(\sqrt{\frac{1}{m}}\right).$
For any $x\in{\bf R}$, the empirical CDFs are defined as:
@@
$D_{i}(x)=\left|\mathbf{1}_{\{\hat{Q}_{i}\leq x\}}-\mathbf{1}_{\{Q_{i}\leq x\}}\right|.$
@@
$|\hat{F}_{m}(x)-F_{m}(x)|\leq\frac{1}{m}\sum_{i=1}^{m}D_{i}(x).$
Since $|\hat{Q}_{i}-Q_{i}|\leq\delta_{m}$, $D_{i}(x)$ can be non-zero only if $Q_{i}\in H(x)$, where $H(x)$ is the interval:
$H(x)=\left\{y\in{\bf R}:|y-x|\leq\delta_{m}\right\}.$
@@
$E\left[D_{i}(x)\right]\leq P\left(Q_{i}\in H(x)\right)\leq C_{\mathrm{up}}\cdot(2\delta_{m})$
@@
$\sup_{x\in{\bf R}}E\left[D_{i}(x)\right]\leq C_{\mathrm{up}}\cdot(2\delta_{m}).$
@@
$E\left[\sup_{x\in{\bf R}}\left|\frac{1}{m}\sum_{i=1}^{m}D_{i}(x)-E[D_{i}(x)]\right|\right]\leq K\sqrt{\frac{1}{m}}.$
@@
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F_{m}(x)|\leq\sup_{x}E[D_{i}(x)]+O_{P}\left(\sqrt{\frac{1}{m}}\right)=O\left(\delta_{m}\right)+O_{P}\left(\sqrt{\frac{p}{m}}\right).$
@@
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\leq O\left(\delta_{m}\right)+O_{P}\left(\sqrt{\frac{1}{m}}\right).$
@@
Under the conditions of Lemma B.1, if $\delta_{m}\to 0$ as $m\to 0$, then the approximate $\hat{F}_{m}$ based on $\hat{Q}_{1},\cdots,\hat{Q}_{m}$ in our problem satisfies that $\hat{F}_{m}$ converges to $F$ uniformly in probability:
$\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\xrightarrow{P}0\quad\text{as }m\to\infty.$
The next tool we need is about a conditional distribution. In particular, suppose a $d$-dimensional random variable $X=(X_{1},\ldots,X_{d})^{\prime}\sim F$, we aim to estimate the conditional CDF
$F^{\,d\mid 1:(d-1)}(x\,\big{|}\,x_{1},\dots,x_{d-1})\;=\;P\!\bigl{(}X_{d}\leq x\,\big{|}\,X_{1}=x_{1},\dots,X_{d-1}=x_{d-1}\bigr{)}.$
This is because we have to estimate the conditional CDF when applying the DIP procedure by following the probability chain rule. In the following result, we will use a kernel estimator. Let $K:{\bf R}^{d-1}\to[0,\infty)$ be a bounded, continuous kernel with $\int_{{\bf R}^{d-1}}K(u)\,du=1$, and let $h>0$ be a bandwidth parameter. Define the kernel-based conditional CDF estimator:
$\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\,\big{|}\,x_{1},\dots,x_{d-1})\;:=\;\frac{\sum_{i=1}^{m}\mathbf{1}\!\bigl{\{}\hat{Q}_{i,d}\leq x\bigr{\}}\;K\!\Bigl{(}\tfrac{x_{1}-\hat{Q}_{i,1}}{h},\dots,\tfrac{x_{d-1}-\hat{Q}_{i,d-1}}{h}\Bigr{)}}{\sum_{i=1}^{m}K\!\Bigl{(}\tfrac{x_{1}-\hat{Q}_{i,1}}{h},\dots,\tfrac{x_{d-1}-\hat{Q}_{i,d-1}}{h}\Bigr{)}}.$

### v3

@@ 7 Examples of real-world data sets
Caltech Facebook network. The first example data set is the Facebook social network between students in California Institute of Technology (Caltech), collected by Traud et al. (2012). Each node in the network is a student and the edges are the Facebook connections between students. We process the data set following Wang and Rohe (2016) and Li et al. (2022) using the 2-core algorithm. The resulting data set contains 734 students with an average degree of 45.29. Similar to our previous experiments, we randomly select half of the students to hold out. The network of the other half of the students ($n=367$) is then privatized and released. The inner product latent space model is used for model fitting, and the latent dimension $d=6$ is selected by the edge cross-validation method of Li et al. (2020a) on the hold-out network. We enforce a strong privacy requirement at the level of $\varepsilon=1$. The distributions of the local statistics in the true and released network are calculated.
@@
Figure 4: The distributions of five local statistics of the privatized Caltech social network with privacy budget $\varepsilon=1$.
Figure 4 displays the resulting distributions of the five local statistics in four networks as introduced in Section 6: the original network (True), the network released GRAND, the network generated using the naive Laplace mechanism (Laplace) and the non-private network from standard model estimation (Hat). It can be seen that the privatized network from GRAND matches the true network well in all of the five metrics. It also substantially outperforms the naive Laplace method, which completely misses the pattern. Meanwhile, GRAND maintains a slightly deviated but similar performance compared to the non-private Hat network. In particular, the harmonic closeness of the original network exhibits a bi-modal pattern, with one tiny lower mode (on the left end of the figure panel). This subtle pattern is also well captured by GRAND.
@@
Figure 5: The distributions of five local properties of the privatized statisticians’ coauthorship network with privacy budget $\varepsilon=1$.
Statisticians’ collaboration network. The second example is about the collaboration network between statisticians based on publication data originally collected by Ji and Jin (2016) and processed in Li et al. (2020b). Each node in this network is a statistician and an edge indicates that the two statisticians coauthored at least one paper during the data collection period. We use the same procedure to process the data set as before, and the resulting network has 509 nodes with an average degree of 4.24. This network is much sparser than the Caltech network, indicating a more difficult model fitting. We still hold out half of the nodes and privatize the network structure of the other half using privacy budget $\varepsilon=1$. The inner product latent space model is used and $d=4$ is selected by cross-validation on the hold-out data.
The distributions of the five local statistics in the resulting networks are shown in Figure 5. The released network from GRAND matches the true network reasonably well, and maintains a performance that is very close to that of the Hat network, with a small deviation due to the incorporation of privacy guarantees. Similar to the Caltech example, the true network also exhibits a bi-modal pattern for the harmonic centrality, but the two modes in this case are much closer to each other, which significantly increases the difficulty in preserving them under the introduced perturbations for privacy. The privatized network from GRAND, though does not perfectly recover the magnitudes, still captures the bi-modal pattern.
Additional evaluations on social network data. We also evaluate GRAND on 107 social networks from Ghasemian et al. (2019) and a privatization task for attributed network, demonstrating similarly competitive performance. These can be found in Appendix H and G.
@@ 8 Discussion
This paper defines node-level differential privacy and introduces a novel privatization mechanism, named GRAND, to achieve node-level differential privacy. To conclude our paper, we want to add more discussions on a few aspects of our work.
Advantages of Perturbation-Based Privacy Protection
The GRAND framework functions primarily as a data-perturbation mechanism. The released network retains a node-level bijective correspondence between the original network and the perturbed one. While our framework could theoretically support synthetic generation—by sampling new latent vectors from the estimated distribution $\hat{F}$ to construct a network of arbitrary size—such an approach yields artificial nodes that lack correspondence to real-world entities. The proposed data-perturbation paradigm exhibit several advantages that are critical in practice.
Extensibility and decentralizability. Real-world data dissemination may involve multiple waves. For instance, the National Longitudinal Study of Adolescent to Adult Health (Add Health) (Harris, 2013) comprises six waves of data collection spanning over 30 years, with new attributes and topological updates continuously appended to the same cohort regularly. A synthetic mechanism is fundamentally incompatible with this longitudinal pattern as nodes generated in previous waves may not match those generated later. In contrast, by maintaining stable node identities, GRAND ensures the private network remains compatible with future data releases and external attribute merging. Another key advantage is that a perturbation mechanism allows for data combination from multiple parties, where each party can separately apply perturbation to their respective datasets. The combined data still follow the DP requirement (with the sum of the privacy budget used by each party) (Dwork et al., 2014).
Robustness to structural over-protection. While node-level DP shields individual topological connections, a utility-maximizing mechanism should preserve meso-scopic features—such as community structure—that characterize groups rather than identifying individuals. A fundamental limitation of synthetic graph mechanisms is the decoupling of structure from identity: the synthetic nodes lose their correspondence to intrinsic ground-truth attributes, rendering original labels (like community membership) meaningless. In contrast, perturbation mechanisms like GRAND operate on the original node set, maintaining the structural alignment required to recover these group-level properties. In Figure 13 of Appendix I, we demonstrate this advantage using the Stochastic Block Model (Holland et al., 1983), a canonical model community structures. We applied spectral clustering to the GRAND-perturbed networks to recover community labels and compared them against the ground truth. GRAND successfully facilitates high recovery accuracy for reasonable $\varepsilon$ regimes. This confirms that GRAND is more robust to the structural over-protection, preserving meso-scopic signals under the privacy constraint, which is inevitable in synthetic approaches.
Future Directions
While laying a foundational cornerstone for solving the node DP problem, this framework also opens numerous avenues for future research. One promising direction is to generalize the proposed method to encode more systematic privacy protection for the hold-out network. As previously discussed, using such hold-out data currently appears unavoidable to achieve the desired privacy guarantees within the current scope of generality. One viable solution is to adopt a $(\varepsilon,\delta)$-differentially private holdout set with theoretical underpinnings and a transfer learning model to apply the learned CDF to the release set. Another potentially useful future direction is to extend the current framework to accommodate more general network models such as the grahpon model (Bickel and Chen, 2009).
@@ Appendix A Proof of Theorem 1
Due to the node-wise estimation procedure, note that the change of any node $i\in[n]$ does not change the estimate of $\hat{Z}_{j}$, $j\neq i$, $j\in[n]$. That means, the estimates $\hat{Z}_{i}$’s are separable. By applying Theorem 2 of Bi and Shen [2023], $\tilde{Z}_{i}$ acquired via Equation (7) is differentially private. Since the generative function $W$ is assumed to be known. We have each node of $\tilde{A}^{11}$ also being differentially private. ∎
@@ Appendix B Proof of Theorem 2 (Marginal Convergence)
The DIP procedure, given $\hat{Z}_{i}$, is to apply univariate DIP transformation by columns. In each step, we use the approximated (conditional) CDF, $\hat{F}_{m}^{j\,|\,1:(j-1)}$ of the $j$-th variable, conditioning on the previous j1j-1 variables, for privatization: for each $i$, we generate $\mathbf{e}_{i}$ from $\text{Laplace}(0,1/\varepsilon)^{d}$ independently, and then compute
$\tilde{Z}_{ij}=(\hat{F}_{m}^{j\,|\,1:(j-1)})^{-1}(G(\hat{F}_{m}^{j\,|\,1:(j-1)}(\hat{Z}_{ij})+e_{ij})).$
@@ B.1 Crucial tools
Let $Q=\{Q_{i}\}_{i=1}^{m}$ be i.i.d. random variables from a distribution $F$ on $\mathbb{R}$, and let $F_{m}(x)$ be the empirical CDF computed from $Q$. Let $\hat{Q}=\{\hat{Q}_{i}\}_{i=1}^{m}$ be a perturbed version of $Q$, which can be potentially dependent. Let $\hat{F}_{m}(x)$ denote the empirical CDF based on the perturbed data $\hat{Q}$.
@@
  • The perturbed data satisfies $\sup_{1\leq i\leq m}\big|\hat{Q}_{i}-Q_{i}\big|\leq\delta_{m}$ for some $\delta_{m}$ that is associated with $m$.
@@
$\sup_{x\in\mathbb{R}}\big|\hat{F}_{m}(x)-F(x)\big|\leq\mathcal{O}_{\mathbb{P}}\left(\delta_{m}\right)+\mathcal{O}_{\mathbb{P}}\biggl(\sqrt{\frac{1}{m}}\biggr).$
@@
$\sup_{x\in\mathbb{R}}\big|\hat{F}_{m}(x)-F(x)\big|\leq\underbrace{\sup_{x\in\mathbb{R}}\big|\hat{F}_{m}(x)-F_{m}(x)\big|}_{\textsf{Perturbation Error}}+\underbrace{\sup_{x\in\mathbb{R}}\big|F_{m}(x)-F(x)\big|}_{\textsf{Sampling Error}}.$
From Glivenko-Cantelli theorem, we have:
$\sup_{x\in\mathbb{R}}\big|F_{m}(x)-F(x)\big|=\mathcal{O}_{\mathbb{P}}\biggl(\sqrt{\frac{1}{m}}\biggr).$
For any $x\in\mathbb{R}$, the empirical CDFs are defined as:
@@
$D_{i}(x)=\big|\mathbf{1}_{\{\hat{Q}_{i}\leq x\}}-\mathbf{1}_{\{Q_{i}\leq x\}}\big|.$
@@
$\big|\hat{F}_{m}(x)-F_{m}(x)\big|\leq\frac{1}{m}\sum_{i=1}^{m}D_{i}(x).$
Since $\big|\hat{Q}_{i}-Q_{i}\big|\leq\delta_{m}$, $D_{i}(x)$ can be non-zero only if $Q_{i}\in H(x)$, where $H(x)$ is the interval:
$H(x)=\left\{y\in\mathbb{R}:|y-x|\leq\delta_{m}\right\}.$
@@
$\mathbb{E}[D_{i}(x)]\leq\mathbb{P}(Q_{i}\in H(x))\leq C_{\mathrm{up}}\cdot(2\delta_{m}),$
@@
$\sup_{x\in\mathbb{R}}\mathbb{E}[D_{i}(x)]\leq C_{\mathrm{up}}\cdot(2\delta_{m}).$
@@
$\mathbb{E}\left[\sup_{x\in\mathbb{R}}\left|\frac{1}{m}\sum_{i=1}^{m}D_{i}(x)-\mathbb{E}[D_{i}(x)]\right|\right]\leq K\sqrt{\frac{1}{m}}.$
@@
$\sup_{x\in\mathbb{R}}\big|\hat{F}_{m}(x)-F_{m}(x)\big|\leq\sup_{x\in\mathbb{R}}\mathbb{E}[D_{i}(x)]+\mathcal{O}_{\mathbb{P}}\biggl(\sqrt{\frac{1}{m}}\biggr)=\mathcal{O}(\delta_{m})+\mathcal{O}_{\mathbb{P}}\biggl(\sqrt{\frac{1}{m}}\biggr).$
@@
$\sup_{x\in\mathbb{R}}\big|\hat{F}_{m}(x)-F(x)\big|\leq\mathcal{O}_{\mathbb{P}}(\delta_{m})+\mathcal{O}_{\mathbb{P}}\biggl(\sqrt{\frac{1}{m}}\biggr).$
@@
Under the conditions of Lemma B.1, if $\delta_{m}\to 0$ as $m\to\infty$, then the approximate $\hat{F}_{m}$ based on $\hat{Q}_{1},\cdots,\hat{Q}_{m}$ in our problem satisfies that $\hat{F}_{m}$ converges to $F$ uniformly in probability:
$\sup_{x\in\mathbb{R}}\big|\hat{F}_{m}(x)-F(x)\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m\to\infty.$
The next tool we need is about a conditional distribution. In particular, suppose a $d$-dimensional random variable $X=(X_{1},\ldots,X_{d})^{\top}\sim F$, we aim to estimate the conditional CDF
$F^{d\,|\,1:(d-1)}(x\,|\,x_{1},\dots,x_{d-1})=\mathbb{P}(X_{d}\leq x\,|\,X_{1}=x_{1},\dots,X_{d-1}=x_{d-1}).$
This is because we have to estimate the conditional CDF when applying the DIP procedure by following the probability chain rule. In the following result, we will use a kernel estimator. Let $K:\mathbb{R}^{d-1}\to[0,\infty)$ be a bounded, continuous kernel with $\int_{\mathbb{R}^{d-1}}K(u)\,\mathrm{d}u=1$, and let $h>0$ be a bandwidth parameter. Define the kernel-based conditional CDF estimator:
$\hat{F}_{m}^{d\,|\,1:(d-1)}(x\,|\,x_{1},\dots,x_{d-1}):=\frac{\sum_{i=1}^{m}\mathbf{1}_{\{\hat{Q}_{i,d}\leq x\}}K\left(\tfrac{x_{1}-\hat{Q}_{i,1}}{h},\dots,\tfrac{x_{d-1}-\hat{Q}_{i,d-1}}{h}\right)}{\sum_{i=1}^{m}K\left(\tfrac{x_{1}-\hat{Q}_{i,1}}{h},\dots,\tfrac{x_{d-1}-\hat{Q}_{i,d-1}}{h}\right)}.$